OpenStudy (anonymous):

Use the following to determine the distance A (13,7) B (-9,4)

OpenStudy (anonymous):

Hers is the formula for your question \[\sqrt{(y2-y1)^{2}+(x2-x1)^{2}}\] \[\sqrt{9+484}\] 21.2 is answer

OpenStudy (anonymous):

1) 5 2) √493 3) √137 4) √112

OpenStudy (anonymous):

Sorry answer is \[\sqrt{493}\]

OpenStudy (anonymous):

\[\sqrt{(x2-x1)^2 + (y2-y1)^2}\] = \[\sqrt{(22^2 + 3^2)}\] = \[\sqrt{484+9}\] = \[\sqrt{493}\]